Fermi National Laboratory
 

Accelerator Update for November 13, 2006 - November 15, 2006

 

Monday November 13

The day shift began with Operations monitoring the Tevatron (TeV) store 5065 (105.31E30) and a Recycler stash (104.3E10), with the Antiproton Source (Pbar) stacking (31.0E10), with Linac, Booster, and the Main Injector (MI) operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the store, and with MiniBooNE, and NuMI taking beam.

At 8:59 AM, LCW technicians at CUB reported that they were switching the MI LCW columns: a column lowers the conductivity of water. Usually this switch isn’t a problem, but a pipe broke. The techs switched back to recirculation mode. Unfortunately, this caused a water pressure pulse that caused the MI-50 LCW pump to trip off, along with several power supplies and MECAR. Operators were able to reset these devices. Still, the MI beam had problems due to MECAR for about 20 minutes.

Operations prepared Pbar for an access at 1:03 PM to search for a LCW leak.

At 1:45 PM, Pbar experts reported finding the LCW leak in AP1, in the transport line, on a quadrupole magnet (IQ19). LCW techs began repairs. This held off beam coming from Booster and going to NuMI, and MiniBooNE until 2:35 PM.

Operations had beam to NuMI and MiniBooNE by 2:40 PM.

At 3:11 PM, a vertical separator spark at TeV sector D17 caused a brief interruption of luminosity.

Pbar resumed stacking around 3:23 PM.

Recycler experts reported at 6:10 PM that they had completed a damper study.

The Recycler completed the fourth and last stash of antiprotons for the day at 10:26 PM.

A Recycler expert started another study at 10:29 PM.

At 11:33 PM, a Booster kicker (MKS06) tripped off. Operators were able to turn the supply on, but turned it off due to internal sparking. Operations called in an expert.

Tuesday November 14

The midnight shift began with Operations monitoring store 5065 (42.01E10), a stash (245.2E10), and a stack (30.7E10), with beam off due to a Booster kicker problem, and with NuMI and MiniBooNE waiting for beam.

The Booster expert completed his kicker repairs. Beam returned at 12:43 AM.

The Recycler completed the second and last stash for the midnight shift around 5:42 AM.

Operations shut off beam to NuMI at 6 AM. NuMI will be off until Thursday while they install a new chilled water compressor.

Experts began a one-hour Tevatron Electron Lens (TEL2) study at 6:33 AM.

Operations terminated store 5065 at 7:33 AM.

Operations put the TeV into shot setup at 7:46 AM.

Operations turned off beam at 9:33 AM and began making MI and Booster ready for access. MI will repair an internal cavity LCW leak on MIRF5. The Run Coordinator will allow other repair and maintenance work during the MI job.

Operations established store 5066 at 9:40 AM with an initial luminosity of 187.86E30.

The MI access began at 10:14 AM.

Booster came out of access at 12:19 PM.

A Booster expert established beam to the Booster dump at 1:03 PM.

MI came out of access at 3:43 PM.

Operations established beam to MiniBooNE at 4:14 PM.

Operations established beam to the MI dump at 4:24 PM.

Pbar began stacking at 5:34 PM.

Operations put the Recycler and Pbar into shot setup at 5:51 PM to stash antiprotons.

The Recycler completed its fifth and last stash for the day at 6:29 PM.

At 6:35 PM, the lithium lens tripped off due to an external interlock. Operators investigated and then called in an expert.

At 7:49 PM, a Booster RF station (BRF19) tripped off. Operators managed to coax the station on ten minutes later.

A Pbar expert reported at 7:56 PM that the lithium lens had tripped off due to a fuse failure.

At 8:38 PM, the 400MeV Area suffered some tuning problems that might have been caused by the Booster chopper (B:CHOP). Operations called in a Booster expert.

The Booster expert reported at 10:16 PM that a high voltage cable for the chopper had popped off.

An expert reported at 11:30 PM that he had replaced the failed fuse and fuse holder for the lithium lens and replaced a switch used to drain energy from the capacitor bank before allowing access.

Pbar resumed stacking at 11:35 PM.

Wednesday November 15

The midnight shift began with Operations monitoring store 5066 (53.18E30) and a stash (56.0E10), with Pbar stacking (27.7E10), with NuMI off for chiller repairs, and with MiniBooNE taking beam.

At 1:56 AM, Operations prepared the Recycler to stash, but an emittance blow up caused them to put the stash on hold and cool down the antiprotons. Operations reported that they lost about 1.5E10 of the stash. An expert was called in.

At 4:20 AM two more Recycler experts came in to work on the damper system.


At 7:43 AM, Operations reported the TeV tilt sensors’ account of the Kuril Islands earthquake (located between Russia’s Kamchatka peninsula and Japan’s Hokkaido Prefecture), which occurred at 11:14 Coordinated Universal Time, 5:14 AM Central Standard time. The earthquake registered a magnitude of 8.3. The TeV colliding beams moved, but there was no quench or loss of luminosity.

For detailed information on the Kuril Islands earthquake please go the USGS Earthquake Hazards Program.

The Plans for Wednesday

The plans for today are to complete the work on the Recycler damper system and stash antiprotons. At noon, store 5066 will be terminated. CDF will be allowed a five-hour access. The accelerator will then resume stacking, stashing, and storing.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/15/2006    email Fermilab